OpenStudy (anonymous):

What is the measure of each interior angle in a regular 15-gon? A. 167° B. 108° C. 156° D. 120°

OpenStudy (anonymous):

@Flvs.net

OpenStudy (flvs.net):

The measure for each interior angle of a regular 15-gon would be "156". Equation: (n-2)180/15 n= the number of sides There is 2,340 degrees in a 15 gon, so you divide that by 15 which equals 156.

OpenStudy (anonymous):

since all of you people are here what do you think

OpenStudy (flvs.net):

156, like it says in my equation.

OpenStudy (mrnood):

the sum of all EXTERIOR angles in any polygon is 360 The interior angle is the complement of that

OpenStudy (hybrik):

To do this you have to do 360/15, that answe will give you each angle

OpenStudy (anonymous):

yea ik. :)

OpenStudy (mrnood):

@hybrik that gives the exterior angle - NOT the interior

OpenStudy (flvs.net):

@MrNood I gave you a medal. :D

OpenStudy (mrnood):

thx

OpenStudy (welshfella):

another way to do this is to use the fact that the sum of all the exterior angles of a polygon add up to 360 so each internal angle in a regular polygon = (180 - (360/n)) where n = number of sides

OpenStudy (flvs.net):

@MrNood Can you give me one back?

OpenStudy (hybrik):

You have to do 2340/15 as Flvs said @MrNood

OpenStudy (welshfella):

oh - thats all been said earlier! - I didn't know!